OpenStudy (anonymous):

Could anyone explain how to take the derivative of 10^(x^2-1)

OpenStudy (anonymous):

so just 10^x as the 2-1 is 1?

OpenStudy (anonymous):

Errr hold on I'll make it more clear.

OpenStudy (anonymous):

\[10^{x ^{2}-1}\]

OpenStudy (turingtest):

do you know how to take the derivative of \(f(x)=a^x\) ?

OpenStudy (anonymous):

you would have to start with the chain rule

OpenStudy (anonymous):

I'm pretty sure I've done something like that before but I'm blanking right now.

OpenStudy (turingtest):

do you recall implicit differentiation?

OpenStudy (anonymous):

Yup.

OpenStudy (turingtest):

\[y=a^x\]\[\ln y=x\ln a\]\[\frac{y'}y=\ln a\]\[y'=y\ln a=a^x\ln a\]

OpenStudy (anonymous):

Ah! I remember now! Thank you, I think I've got it now.

OpenStudy (turingtest):

welcome, and don't forget to use the chain rule too :)